# Cold day-side winds shape large leading streams in evaporating exoplanet atmospheres

### v1

@@
Cold day-side winds shape large leading streams in evaporating exoplanet atmospheres
Recent observations of planetary atmospheres in HAT-P-32 b and HAT-P-67 b reveal extensive outflows reaching up to hundreds of planetary radii. The helium 1083 nm light curves for these planets, captured across their full orbits, show notable asymmetries: both planets display more pronounced pre-transit than post-transit absorptions, with HAT-P-67 b being the more extreme case of that geometry. Using three-dimensional (3D) hydrodynamic simulations, we identify key factors influencing the formation of a dense leading outflow stream and characterize its morphology. Our models suggest that such a geometry of escaped material is caused by a relatively cold outflow of high mass-loss rate, launched preferentially from the planet’s day side. From the simulations we calculate synthetic He i 1083 nm spectra that show large absorption depths and irregular line profiles due to complex gas kinematics. We find that the measurements of the He i 1083 nm equivalent width and the velocity shift relative to the planet’s rest frame, observed over a significant portion of the planet’s orbital phase, can provide important constraints on the outflow properties and its interaction with the stellar wind.
@@
The study of atmospheric escape in planets is crucial for understanding the mass loss and long-term evolution of planetary atmospheres (e.g., Owen, 2019). Spectroscopic observations, particularly in the He i 1083 nm line, have become essential tools for tracing the kinematics and structures of these outflows (e.g., Oklopčić & Hirata, 2018; Spake et al., 2018; Nortmann et al., 2018; Allart et al., 2018; Salz et al., 2018). Recent He i 1083 nm transit observations conducted with the Habitable-zone Planet Finder Spectrograph (HPF) on the 10-meter Hobby-Eberly Telescope (HET) for HAT-P-32 b (Zhang et al., 2023) and HAT-P-67 b (Gully-Santiago et al., 2023) have revealed planetary outflows extending over hundreds of planetary radii. HAT-P-67 b is particularly intriguing due to its exceptionally extended pre-transit outflow, reaching orbital phases as early as , followed by a sharp drop immediately after the optical transit. The helium light curve of HAT-P-32 b also shows a notable asymmetry, with a stronger pre-transit phase that lasts an hour longer than the post-transit phase. Both planets display remarkably high helium excess depths, with 8.2% for HAT-P-32 b and 10% for HAT-P-67 b. These predominantly leading, high-density outflows differ from the more typical scenario in which the outflow is more extended on the trailing side, behind the planet (e.g., Ehrenreich et al., 2015; Nortmann et al., 2018; Guilluy et al., 2020; Spake et al., 2021; Ben-Jaffel et al., 2022; Tyler et al., 2024). This unusual morphology motivates us to investigate what outflow properties can cause such a configuration, using 3D hydrodynamic simulations.
@@
The hot Saturn HAT-P-67 b (Zhou et al., 2017; Gully-Santiago et al., 2023) orbits an evolved F5 subgiant ( K) at a distance of au, with an orbital period of days and a transit duration of 7.0 hours. The planet’s proximity to its host star, coupled with the stellar evolution through the subgiant phase and increased luminosity, results in notably high insolation for HAT-P-67 b. The planet has an equilibrium temperature of K, a reported upper mass limit of , and a radius of , making it inflated with a low surface gravity of dex. HAT-P-67 b is assumed to be on a circular orbit, and Doppler tomographic observations have shown that the planet’s orbital plane is closely aligned with the equatorial plane of its host star (Zhou et al., 2017; Sicilia et al., 2024). The HET observations suggest a high planetary mass-loss rate on the order of g s-1 based on the comparison with spherically symmetric (1D) Parker wind models. Additional high-resolution transit data were obtained by the CARMENES spectrograph (Bello-Arufe et al., 2023) and the combined spectrographs GIARPS (GIANO-B + HARPS-N) (Sicilia et al., 2024). While the short transit baselines limited the capture of the outflow’s full extent, both studies noted a helium redshift during the transit. Notably, Fig. 8 of (Bello-Arufe et al., 2023) shows greater pre-transit absorption depth, consistent with Gully-Santiago et al. (2023).
@@
The hot Jupiter HAT-P-32 b (Hartman et al., 2011; Wang et al., 2019) orbits a late-F to early-G star ( K) at a distance of au, with an orbital period of days and equilibrium temperature of K. The transit duration is 3.12 hours and the orbit has eccentricity of . The planet has a mass of and a radius of , leading to a surface gravity of dex.
@@
The characteristics of HAT-P-32 b and HAT-P-67 b have many similarities. Due to their small orbital separation and the high mass ratio between the star and the planet, both planets are close to filling their Roche lobe, with for HAT-P-32 b and for HAT-P-67 b. CARMENES observations by Czesla et al. (2022) detected hydrogen (H) and helium (He-I) outflows from HAT-P-32 b. The spectral lines show irregular shapes and significant, time-dependent variations. 1D hydrodynamic modeling of these data by Yan et al. (2024) suggests that the planet’s atmosphere has solar metallicity. 1D Parker wind models by Lampón et al. (2023) and 3D hydrodynamic simulations by Zhang et al. (2023) estimate a mass-loss rate of approximately g s-1, similar to the estimate for HAT-P-67 b.
@@
This paper aims to model large-scale structures of asymmetric outflows as observed in HAT-P-67 b and HAT-P-32 b, with a particular focus on the conditions under which a high-density leading stream emerges. It is organized as follows: In Section 2, we introduce the simulation setup and approach. Section 3 addresses the outflow conditions that lead to a leading tail only and explores the impact of the stellar wind on the outflow. Section 4 places our results in context and discusses their broader implications. Finally, we present our conclusions in Section 5.
@@ 2 Methods
To simulate the planetary outflow and its interaction with the stellar wind and the gravitational field of the host star, we have used three-dimensional (3D) hydrodynamical models and post-processed them to estimate the effects of the escaping planetary material on the transit spectra, focusing on the metastable helium triplet around nm. The primary goal of this study is to identify the parameters responsible for the presence of a leading outflow stream, characterize the structural features of this planetary wind morphology, and determine how these features are reflected in helium transit observations. The simulations presented in this paper are modeled on parameters similar to those of HAT-P-67 b.
@@ 2.1 Simulations
We simulate a planet orbiting its host star, incorporating thermal winds from both the planet and the star, using the 3D Eulerian (magneto)hydrodynamic code Athena++111https://github.com/PrincetonUniversity/athena, version 2021 (Stone et al., 2020). For these hydrodynamic simulations, we use the setup previously developed in MacLeod & Oklopčić (2022) and extended to account for planetary day-night anisotropies in Nail et al. (2024). In the following, we outline the key features of our simulations and the modifications made for this study, building on the methods described in Section 2 of those works.
@@
The simulation solves the equations for mass, momentum, and energy conservation in an inviscid gas. We use the ideal gas equation of state with an almost isothermal adiabatic index of , which results in nearly isothermal behavior along adiabats. However, temperatures at the points where the planetary winds originate from the surface can vary, allowing us to incorporate temperature differences between the planet’s day side and night side, as described by Nail et al. (2024). We consider the gravitational influence of the planet and the star, while neglecting the back-reaction of the outflow on the planetary orbit and the impact of 1083 nm radiation pressure on the gas dynamics.
@@
We insert hydrodynamic winds into our simulations by parameterizing the conditions at the bases of the stellar and planetary winds. The pressure on the stellar surface is determined by the constant surface density and the hydrodynamic escape parameter , which represents the ratio of gravitational potential energy to thermal energy . For all models, we use and g cm-3. Similarly, the pressure at the planetary surface is given by
@@
where represents the hydrodynamic escape parameter for the planet. In models that account for day-night anisotropy, the surface pressure can vary across the planetary surface.
@@
Using a spherical polar mesh with the star at the center of the co-rotating frame aligned with the planet’s orbit, our computational domain extends from the star’s surface to a distance of 12.3 au in all directions. The planet is located on the negative x-axis at a distance of au from the star. Both objects rotate with the orbital frequency of the planet, , with their angular momenta aligned in the positive z-direction. The base mesh is composed of meshblocks, each containing zones, logarithmically spaced in the direction and evenly spaced in and directions, maintaining nearly cubic zone shapes. We reduce the effective zones in the direction near the poles by averaging conserved quantities to avoid extreme aspect ratios.
@@
To assess the sensitivity of our results to numerical parameters, we varied the spatial resolution with three refinement levels around the planet (increasing by a factor of two) and three levels in the outflow torus. We found that stellar and planetary outflow rates stabilize after a few planetary orbits and converge above the lowest resolution levels (). However, spectral properties, particularly in the extended tails, require higher resolution. Thus, we use for the region around the planet and for the tail in our fiducial settings.
@@
We investigate six models with varying planetary outflow temperatures and geometries, as well as stellar mass-loss rates. Detailed input parameters are provided in Table 1. Models designated with a prefix ”B” have a low value of , which corresponds to a hot planetary wind. Models designated with a prefix ”S” employ a high , resulting in a cold wind. Models B1 and S1 assume an isotropic planetary outflow, whereas all other models incorporate a day-side dominant outflow configuration using the anisotropic planetary wind model introduced in our previous work Nail et al. (2024), detailed in Section 2.2. In these anisotropic scenarios, the pressure on the night side of the planet is only 1% of that on the day side, as described by the parameter . Models S3 and S4 are similar to the anisotropic cold model S2, but we gradually increased the stellar mass-loss rate. The system parameters, adopted from HAT-P-67 b, are identical for all models: au = cm, g, cm, g cm-3, g, cm.
| model | |||||
| [%] | [K] | [km s-1] | [km s-1] | ||
| B1 | 3 | 100 | 14 100 | 9.7 | 9.7 |
| B2 | 3 | 1 | 14 100 | 9.7 | 1.0 |
| S1 | 9 | 100 | 4 700 | 5.6 | 5.6 |
| S2 | 9 | 1 | 4 700 | 5.6 | 0.6 |
| S3 | 9 | 1 | 4 700 | 5.6 | 0.6 |
@@
Notes. We present two hot wind models (B) and four cold wind models (S). For each, is the hydrodynamic escape parameter, and is the pressure fraction at the antistellar point relative to the substellar point. The table lists the substellar temperature , calculated with (although this varies as a function of the outflow’s location); and the sound speed at both substellar (sub) and antistellar (anti) points on the planetary surface, which has a constant density . Models S3 and S4, have a higher stellar mass-loss rate compared to model S2.
@@ 2.2 Radiative transfer
In our radiative transfer analysis, we adopt the approach outlined by MacLeod & Oklopčić (2022). We assume a solar gas composition with mass fractions of hydrogen, helium, and metals represented by X = 0.738, Y = 0.248, Z = 0.014, respectively, throughout the computational domain. We assume detailed equilibrium in each simulation grid cell, treating the system as being in a steady state. To compute the unattenuated photoionization rate, , we adopted an intermediate flux between the low and high XUV synthetic spectra of HAT-P-67 as presented in Gully-Santiago et al. (2023) (see their Fig. 12). Combined with the optical depth , this yields the photoionization rate . The optical depth to hydrogen-ionizing photons is iteratively calculated for each cell in the simulation grid.
To generate synthetic helium spectra, we compute the integrated optical depth along rays extending from the stellar surface to the observer in a simulation snapshot. A Voigt line profile is imposed, incorporating a Gaussian component dependent on the local gas temperature. Rays crossing the planet’s interior are excluded, and the contribution of rays is weighted according to the stellar quadratic limb darkening law. An impact parameter of 0.5, as determined by Gully-Santiago et al. (2023), is used in the calculations of the spectra. Rotational effects of the stellar disk are not considered.
Due to the larger optical depths in our simulations, more iterations are required to determine the ion populations compared to MacLeod & Oklopčić (2022) (see their equations 9 and 10). The number of required iterations increases with the up-scaling of the density in the snapshot. Therefore, we use an iterative scheme to ensure convergence of level populations even when the optical depth is of order unity. We uploaded the updated radiative transfer code together with simulation material on Zenodo22210.5281/zenodo.13988501.
@@ 3.1 Formation of a predominantly leading stream
In this section, we systematically explore a range of parameters in our simulations to generate a high-density leading outflow stream. Our analysis is based on a comparison with the He i 1083 nm observations of HAT-P-67 b and HAT-P-32 b, shown in Fig. 1. The top panels display the total He i 1083 nm equivalent width, including the stellar line and the excess absorption from the planet, as a function of the planet’s orbital phase. Notably, the pre-transit absorption, likely caused by a high-density leading stream, persist longer than the post-transit absorption in both planets. In the bottom panel, we present the observed velocity shifts of the residual planetary helium line. The line is significantly blueshifted during pre-transit observations. While the in-transit observations seem to be variable and partly follow the planetary movement, the post-transit spectra do not. Instead, they either remain in the stellar rest frame or shift slightly towards shorter wavelengths in the later phases.
To understand the factors influencing the formation of an extended leading outflow, we compare and analyze models B1, B2, S1 and S2. In Fig. 2, we present their outflow morphologies, illustrating the metastable helium number density of the planetary wind in the orbital mid-plane as calculated from the radiative transfer analysis. In Fig. 3, we display the corresponding light curves, showing the equivalent width of the helium line as a function of orbital phase. Unlike the light curves shown at the top of Fig. 1, here we display only the planet’s excess absorption, with the stellar contribution excluded.
Due to the scale-invariance of hydrodynamic equations, we can uniformly adjust the pressure and density of a simulation snapshot during post-processing without altering the outflow morphologies. This allows us to modify the planetary and stellar mass-loss rates, without changing their ratio, before generating the synthetic spectra. We used this capability to match the model light curve to the mid-transit absorption of HAT-P-67 b. We provide the resulting mass-loss rates of these models in Table 2.
@@
Firstly, we examined model B1, which features a hot planetary isotropic outflow (Fig. 2, top left). In this configuration, the strong planetary wind creates a cavity against the stellar wind pressure. Due to the high temperature of the outflow, the radial velocity component dominates over the shear velocity, resulting in a bubble-shaped morphology. Optical depth contours of metastable helium reveal a symmetric density distribution around the planet. Beyond the cavity, both leading and trailing outflows emerge. However, the density in the extended outflow is much lower than in the vicinity of the planet. This is reflected in the light curve, shown by the red line in Fig. 3.
@@
In an attempt to produce an asymmetric light curve with higher pre-transit absorption, we implemented a day-side dominated outflow compared to the first model. This second model, B2 (Fig. 2, top right), is similar to model B1 but features a slightly more extended cavity on the day side. As discussed in Nail et al. (2024), a net day-to-night side flow arises due to the pressure gradient from the substellar to antistellar point, resulting in increased density on the trailing side. Additionally, the pressure in the leading arm is insufficient to counteract the stellar wind. Consequently, the gas is pushed to the trailing side of the planet, contributing to the extended trailing arm. However, the density in the trailing side is too low to contribute significantly to post-transit observations (Fig. 3, green line).
@@
A distinct morphology emerges in the models S1 and S2, featuring a colder planetary wind, corresponding to a high hydrodynamic escape parameter (). Unlike the morphologies previously described, high-density streams surround the planet (Fig. 2, bottom). The densities in the immediate vicinity of the planet are so high that a portion of the XUV radiation is blocked, resulting in a low ionization fraction and, consequently, a low population of metastable helium. The day side outflow launched toward the star, crosses the inner Lagrange point, and is drawn into a Keplerian orbit after overcoming the gravitational potential of the planet. The outflow is accelerating in the positive x-direction, extending potentially to hundreds of planetary radii. On the night side of the planet, the outflow passes through the Lagrange point and undergoes similar entrainment into a Keplerian orbit, accelerating in the negative x-direction. Since the point is further outward relative to the orbit than , the trailing gas moves slower, while the gas leading the planet moves faster.
@@
While the isotropic model S1 produces a more symmetric outflow surrounding the planet, resulting in a toroidal structure around the star, model S2 demonstrates a higher density in the leading stream. Model S2 is the only configuration that displays significantly greater pre-transit absorption compared to post-transit absorption (see the blue light curve in Fig. 3). In conclusion, our findings suggest that a leading tail extending over hundreds of planetary radii is achieved when a cold planetary wind originates exclusively from the day side. Simulation model S2 seems to best reproduce this behavior; however, the pre-transit absorption seems to be underestimated compared to the observational results shown in the top panel of Fig. 1.

### v2

# Cold dayside winds shape large leading streams in evaporating exoplanet atmospheres

@@
Cold dayside winds shape large leading streams in evaporating exoplanet atmospheres
Recent observations of planetary atmospheres in HAT-P-32 b and HAT-P-67 b reveal extensive outflows reaching up to hundreds of planetary radii. The helium 1083 nm light curves for these planets, captured across their full orbits, show notable asymmetries: both planets display more pronounced pre-transit than post-transit absorptions, with HAT-P-67 b being the more extreme case. Using 3D hydrodynamic simulations, we identified the key factors influencing the formation of a dense leading outflow stream and characterized its morphology. Our models suggest that such a geometry of escaped material is caused by a relatively cold outflow with a high mass-loss rate, launched preferentially from the planet’s dayside. From the simulations we calculated synthetic He i 1083 nm spectra that show large absorption depths and irregular line profiles due to complex gas kinematics. We find that the measurements of the He i 1083 nm equivalent width and the velocity shift relative to the planet’s rest frame, observed over a significant portion of the planet’s orbital phase, can provide important constraints on the outflow properties and its interaction with the stellar wind.
@@
The study of atmospheric escape in planets is crucial for understanding the mass loss and long-term evolution of planetary atmospheres (e.g., Owen, 2019). Spectroscopic observations, particularly in the He i 1083 nm line, have become essential tools for tracing the kinematics and structures of these outflows (e.g., Oklopčić & Hirata, 2018; Spake et al., 2018; Nortmann et al., 2018; Allart et al., 2018; Salz et al., 2018). Recent He i 1083 nm transit observations conducted with the Habitable-zone Planet Finder Spectrograph (HPF) on the 10-meter Hobby-Eberly Telescope (HET) of HAT-P-32 b (Zhang et al., 2023) and HAT-P-67 b (Gully-Santiago et al., 2024) have revealed planetary outflows extending over hundreds of planetary radii. HAT-P-67 b is particularly intriguing due to its exceptionally extended pre-transit outflow, reaching orbital phases as early as , which is followed by a sharp drop immediately after the optical transit. The helium light curve of HAT-P-32 b also shows a notable asymmetry, with a stronger pre-transit phase that lasts an hour longer than the post-transit phase. Both planets display remarkably high helium excess depths: 8.2% for HAT-P-32 b and 10% for HAT-P-67 b. These predominantly leading, high-density outflows differ from the more typical scenario in which the outflow is more extended on the trailing side, behind the planet (e.g., Ehrenreich et al., 2015; Nortmann et al., 2018; Guilluy et al., 2020; Spake et al., 2021; Ben-Jaffel et al., 2022; Tyler et al., 2024). This unusual morphology motivated us to investigate what outflow properties can cause such a configuration using 3D hydrodynamic simulations.
@@
The hot Saturn HAT-P-67 b (Zhou et al., 2017; Gully-Santiago et al., 2024) orbits an evolved F5 subgiant (effective temperature K) at a distance of au, with an orbital period of days and a transit duration of 7.0 hours. The planet’s proximity to its host star, coupled with the stellar evolution through the subgiant phase and increased luminosity, results in notably high insolation for HAT-P-67 b. The planet has an equilibrium temperature of K, a reported upper mass limit of , and a radius of , making it inflated with a low surface gravity of dex. HAT-P-67 b is assumed to be on a circular orbit, and Doppler tomographic observations have shown that the planet’s orbital plane is closely aligned with the equatorial plane of its host star (Zhou et al., 2017; Sicilia et al., 2024). The HET observations suggest a high planetary mass-loss rate on the order of g s-1 based on the comparison with spherically symmetric (1D) Parker wind models (Gully-Santiago et al., 2024). Additional high-resolution transit data were obtained by the CARMENES111Calar Alto high-Resolution search for M dwarfs with Exoearths with Near-infrared and optical Échelle Spectrographs. spectrograph (Bello-Arufe et al., 2023) and the combined spectrograph GIARPS (GIANO-B + HARPS-N; Sicilia et al., 2024). While the short transit baselines limited the capture of the outflow’s full extent, both studies (Bello-Arufe et al., 2023; Gully-Santiago et al., 2024) noted a helium redshift during the transit. Notably, Fig. 8 of Bello-Arufe et al. (2023) shows greater pre-transit absorption depth, consistent with Gully-Santiago et al. (2024).
@@
The hot Jupiter HAT-P-32 b (Hartman et al., 2011; Wang et al., 2019) orbits a late-F to early-G star ( K) at a distance of au, with an orbital period of days and an equilibrium temperature of K. The transit duration is 3.12 hours, and the orbit has an eccentricity of . The planet has a mass of and a radius of , resulting in a surface gravity of dex.
@@
The characteristics of HAT-P-32 b and HAT-P-67 b have many similarities. Due to their small orbital separation and the high mass ratio between the star and the planet, both planets are close to filling their Roche lobe, with for HAT-P-32 b and for HAT-P-67 b. CARMENES observations by Czesla et al. (2022) detected hydrogen (H) and helium (He-I) outflows from HAT-P-32 b. The spectral lines show irregular shapes and significant, time-dependent variations. 1D hydrodynamic modeling of these data by Yan et al. (2024) suggests that the planet’s atmosphere has a solar metallicity. 1D Parker wind models by Lampón et al. (2023) and 3D hydrodynamic simulations by Zhang et al. (2023) estimate a mass-loss rate of approximately g s-1, similar to the estimate for HAT-P-67 b.
This paper aims to model large-scale structures of asymmetric outflows as observed in HAT-P-67 b and HAT-P-32 b, with a particular focus on the conditions under which a high-density leading stream emerges. Our work tries to understand these systems’ qualitative features rather than quantitatively fitting their observations.
@@
In Sect. 2 we introduce the simulation setup and approach. Section 3 addresses the outflow conditions that lead to a leading tail only and explores the impact of the stellar wind on the outflow. Section 4 places our results in context and discusses their broader implications. Finally, we present our conclusions in Sect. 5.
@@ 2 Methods
To simulate the planetary outflow and its interaction with the stellar wind and the gravitational field of the host star, we used 3D hydrodynamical models and post-processed them to estimate the effects of the escaping planetary material on the transit spectra, focusing on the metastable helium triplet around nm. The primary goal of this study is to identify the parameters responsible for the presence of a leading outflow stream, characterize the structural features of this planetary wind morphology, and determine how these features are reflected in helium transit observations. The simulations presented in this paper are modeled on parameters similar to those of HAT-P-67 b.
@@ 2.1 Simulations
We simulated a planet orbiting its host star, incorporating thermal winds from both the planet and the star using the 3D Eulerian (magneto)hydrodynamic code Athena++222https://github.com/PrincetonUniversity/athena, version 2021 (Stone et al., 2020). For these hydrodynamic simulations, we used the setup previously developed in MacLeod & Oklopčić (2022) and extended to account for planetary day-night anisotropies in Nail et al. (2024). In the following, we outline the key features of our simulations and the modifications made for this study, building on the methods described in Sect. 2 of those works.
@@
The simulation solves the equations for mass, momentum, and energy conservation in an inviscid gas. We used the ideal gas equation of state with an almost isothermal adiabatic index of , which results in nearly isothermal behavior along adiabats. However, temperatures at the points where the planetary winds originate from the surface can vary, allowing us to incorporate temperature differences between the planet’s dayside and nightside, as described by Nail et al. (2024). We considered the gravitational influence of the planet and the star, while neglecting the back-reaction of the outflow on the planetary orbit and the impact of 1083 nm radiation pressure on the gas dynamics.
@@
We inserted hydrodynamic winds into our simulations by parameterizing the conditions at the bases of the stellar and planetary winds. The pressure on the stellar surface is determined by the constant surface density, , and the hydrodynamic escape parameter, , which represents the ratio of gravitational potential energy, , to thermal energy, . Similarly, the pressure at the planetary surface is given by
@@
where represents the hydrodynamic escape parameter for the planet. The surface densities of the planet () and the star () remain constant within each model and are provided in Table 1. In models that account for day-night anisotropy, the surface pressure can vary across the planetary surface.
@@
Using a spherical polar mesh with the star at the center of the corotating frame aligned with the planet’s orbit, our computational domain extends from the star’s surface to a distance of 12.3 au in all directions. The planet is located on the negative x-axis at a distance of au from the star. Both objects rotate with the orbital frequency of the planet, , with their angular momenta aligned in the positive z direction. The base mesh is composed of mesh blocks, each containing zones, logarithmically spaced in the direction and evenly spaced in and directions, maintaining nearly cubic zone shapes. We reduced the effective zones in the direction near the poles by averaging conserved quantities to avoid extreme aspect ratios.
@@
To assess the sensitivity of our results to numerical parameters, we varied the spatial resolution with three refinement levels around the planet (increasing by a factor of two) and three levels in the outflow torus. We find that stellar and planetary outflow rates stabilize after a few planetary orbits and converge above the lowest resolution levels (). However, spectral properties, particularly in the extended tails, require higher resolution. Thus, we used for the region around the planet and for the tail in our fiducial settings.
We investigated six models with varying planetary outflow temperatures and geometries, as well as stellar mass-loss rates. Detailed input parameters are provided in Table 1. Models designated with a prefix ”B” have a low value of , (assuming a gas temperature of 14000 K (based on 1D model results from Gully-Santiago et al., 2024) and a mean molecular weight of 0.6, consistent with system properties and uncertainties. However, our simulations indicate that helium line broadening can be kinematic rather than purely thermal, motivating the exploration of cooler winds. To test this, we introduced models, prefixed with ”S,” that assume a colder planetary wind with , corresponding to a temperature of approximately 5000 K, though it could theoretically be even lower.
@@
Models B1 and S1 assume an isotropic planetary outflow, whereas all other models incorporate a dayside dominant outflow configuration using the anisotropic planetary wind model introduced in our previous work Nail et al. (2024), detailed in Sect. 2.2. In these anisotropic scenarios, the pressure on the nightside of the planet is only 1% of that on the dayside, as described by the parameter . Models S3 and S4 are similar to the anisotropic cold model S2, but we gradually increased the stellar mass-loss rate. The system parameters, adopted from HAT-P-67 b, are identical for all models: au = cm, g, cm, g, and cm.
| model | |||||||
| [g cm-3] | [g cm-3] | [%] | [K] | [km s-1] | [km s-1] | ||
| B1 | 3 | 100 | 14 100 | 9.7 | 9.7 | ||
| B2 | 3 | 1 | 14 100 | 9.7 | 1.0 | ||
| S1 | 9 | 100 | 4 700 | 5.6 | 5.6 | ||
| S2 | 9 | 1 | 4 700 | 5.6 | 0.6 | ||
| S3 | 9 | 1 | 4 700 | 5.6 | 0.6 | ||
@@
Notes. We present two hot wind models (B) and four cold wind models (S). For each, is the hydrodynamic escape parameter, and is the pressure fraction at the anti-stellar point relative to the substellar point. The table lists the planetary and stellar surface density ( and , respectively); the substellar temperature , calculated with (although this varies as a function of the outflow’s location); and the sound speed at both substellar (sub) and anti-stellar (anti) points on the planetary surface, which has a constant density . Models S3 and S4 have a higher stellar mass-loss rate compared to model S2.
@@ 2.2 Radiative transfer
In our radiative transfer analysis, we adopted the approach outlined by MacLeod & Oklopčić (2022). We assumed a solar gas composition with mass fractions of hydrogen, helium, and metals represented by X = 0.738, Y = 0.248, and Z = 0.014, respectively, throughout the computational domain. We assumed a detailed equilibrium in each simulation grid cell, treating the system as being in a steady state. To compute the unattenuated photoionization rate, , we adopted an intermediate flux between the low and high extreme-ultraviolet (XUV) synthetic spectra of HAT-P-67 as presented in Gully-Santiago et al. (2024, see their Fig. 12). Combined with the optical depth , this yields the photoionization rate . The optical depth to hydrogen-ionizing photons is iteratively calculated for each cell in the simulation grid.
To generate synthetic helium spectra, we computed the integrated optical depth along rays extending from the stellar surface to the observer in a simulation snapshot. A Voigt line profile is imposed, incorporating a Gaussian component that is dependent on the local gas temperature. Rays crossing the planet’s interior are excluded, and the contribution of rays is weighted according to the stellar quadratic limb darkening law. An impact parameter of 0.5, as determined by Gully-Santiago et al. (2024), is used in the calculations of the spectra. Rotational effects of the stellar disk are not considered.
Due to the larger optical depths in our simulations, more iterations are required to determine the ion populations compared to MacLeod & Oklopčić (2022, see their Eqs. 9 and 10). The number of required iterations increases with the up-scaling of the density in the snapshot. Therefore, we used an iterative scheme to ensure convergence of level populations even when the optical depth is of order unity. The code updates the densities at each iteration based on radiative transfer calculations, including photoionization and recombination processes. The solution is updated until the convergence criterion is met, specifically when the relative change in the ionized hydrogen density, becomes smaller than a predefined threshold (here 0.02%).
@@ 3.1 Formation of a predominantly leading stream
In this section we systematically explore a range of parameters in our simulations to generate a high-density leading outflow stream. Our analysis is based on a comparison with the He i 1083 nm observations of HAT-P-67 b and HAT-P-32 b, shown in Fig. 1. Typically, planetary excess absorption is isolated by calculating the difference between the observed absorption during transit and the baseline stellar absorption observed out of transit. The residual line refers to the observed absorption that remains after accounting for the baseline. This method separates the contribution from the planet’s atmosphere, enabling the measurement of absorption by different atmospheric species. However, if the baseline observed out of transit is contaminated, it could lead to a misinterpretation of the planetary absorption. Therefore, the top panels of Fig. 1 present the combined He i 1083 nm absorption equivalent width (EW) from both the star and the planet. These plots demonstrate that the planetary excess absorption persists even after the transit.
Notably, the pre-transit absorption, likely caused by a high-density leading stream, persist longer than the post-transit absorption in both planets. In the bottom panel, we present the observed velocity shifts of the residual planetary helium line. The velocity shift of the residual helium line refers to the displacement of the observed absorption feature from the expected rest wavelength of the helium line, which indicates the motion of the planetary atmosphere. A blueshift means that the gas is moving toward the observer, while a redshift indicates that the gas is moving away. The residual helium line, of HAT-P-67 b and HAT-P-32 b, is significantly blueshifted during pre-transit observations. While the in-transit observations seem to be variable and partly follow the planetary movement, the post-transit spectra do not. Instead, they either remain in the stellar rest frame or shift slightly toward shorter wavelengths in the later phases.
To understand the factors influencing the formation of an extended leading outflow, we compared and analyzed models B1, B2, S1, and S2. In Fig. 2 we present their outflow morphologies, illustrating the metastable helium number density of the planetary wind in the orbital midplane as calculated from the radiative transfer analysis. In Fig. 3 we display the corresponding light curves, showing the EW of the helium line as a function of orbital phase. To calculate the EW, we integrated the total flux of the helium triplet over the wavelength range of 10827.00 to 10831.75 . Unlike the light curves shown at the top of Fig. 1, here we display only the planet’s excess absorption, with the stellar contribution excluded.
Due to the scale-invariance of hydrodynamic equations, we can uniformly adjust the pressure and density of a simulation snapshot during post-processing without altering the outflow morphologies. This allows us to modify the planetary and stellar mass-loss rates, without changing their ratio, before generating the synthetic spectra. We used this capability to match the model light curve to the mid-transit absorption of HAT-P-67 b. We provide the resulting mass-loss rates of these models in Table 2. Mass-loss rates are determined by tracking the mass changes of the planet and star over time and accounting for mass flux through the simulation boundaries. This method provides a net mass-loss rate for each object, ensuring mass conservation while accounting for mass ejected, accreted, or redistributed within the system. We focused on the steady-state regime by averaging the rates over times greater than s, scaling the results by the predefined scaling factor for each simulation.
@@
Firstly, we examined model B1, which features a hot planetary isotropic outflow (Fig. 2, top left). In this configuration, the strong planetary wind creates a cavity against the stellar wind pressure. Due to the high temperature of the outflow, K at the Hill radius of the planet, the radial velocity component km s-1dominates over the shear velocity km s-1, resulting in a bubble-shaped morphology. Optical depth contours of metastable helium reveal a symmetric density distribution around the axis and the shadow of the planet, in the immediate vicinity of the planet, inside the unshocked cavity. Beyond the cavity, both leading and trailing outflows emerge. However, the density in the extended outflow is much lower than in the vicinity of the planet. This is reflected in the light curve, shown by the red line in Fig. 3.
@@
In an attempt to produce an asymmetric light curve with higher pre-transit absorption, we implemented a dayside-dominated outflow compared to the first model. This second model, B2 (Fig. 2, top right), is similar to model B1 but features a more extended cavity on the dayside compared to the nightside. As discussed in Nail et al. (2024), a net day-to-nightside flow arises due to the pressure gradient from the substellar to anti-stellar point, resulting in increased density on the trailing side. Additionally, the pressure in the leading arm is insufficient to counteract the stellar wind. Consequently, the gas is pushed to the trailing side of the planet, contributing to the extended trailing arm. However, the density in the trailing side is too low to contribute significantly to post-transit observations (Fig. 3, green line).
@@
A distinct morphology emerges in models S1 and S2, featuring a colder planetary wind, corresponding to a high hydrodynamic escape parameter (). Unlike the morphologies previously described, high-density streams surround the planet (Fig. 2, bottom). The densities in the immediate vicinity of the planet are so high that a portion of the XUV radiation is blocked, resulting in a low ionization fraction and, consequently, a low population of metastable helium. The dayside outflow launched toward the star, with km s-1and K, crosses the inner Lagrange point, and is drawn into a Keplerian orbit after overcoming the gravitational potential of the planet. The outflow can potentially extend over hundreds of planetary radii along the planet’s orbit. In the case of a hot wind, the radial velocity component is sufficiently strong to prevent the gas from being pulled back into orbit after passing the point. On the nightside of the planet In the cold wind scenario, the outflow accelerates in the negative x direction, passes through the Lagrange point, and undergoes similar entrainment into a Keplerian orbit. Since the point is further outward relative to the orbit than , the trailing gas moves more slowly, while the gas leading the planet moves faster.
@@
While the isotropic model S1 produces a more symmetric outflow surrounding the planet, resulting in a toroidal structure around the star, model S2 demonstrates a higher density in the leading stream. Model S2 is the only configuration that displays significantly greater pre-transit absorption compared to post-transit absorption (see the blue light curve in Fig. 3). Notably, the helium EW ratio between orbital phases and is highest for model S2, with a value of 9.8, while model S1 shows a significantly lower ratio of 1.9. When comparing this to the observational results for HAT-P-67 b (see Fig.1), and assuming, with some uncertainty, that the planet’s excess absorption can be approximated by subtracting a baseline value of 0.2 from the combined stellar and planetary helium EW light curve, the resulting ratio is approximately 7, which is closer to the value predicted by model S2. In conclusion, our findings suggest that a leading tail extending over hundreds of planetary radii is achieved when a cold planetary wind originates exclusively from the dayside. Simulation model S2 seems to best reproduce this behavior; however, the pre-transit absorption seems to be underestimated compared to the observational results shown in the top panel of Fig. 1.